For principal quantum number of n = 4, the total number of orbitals having  = 3 is ?
 

Grade:12

4 Answers

Arun
6 years ago
Dear Student
 
Answer is 7.
 
m = -3, -2, -1, 0, 1, 2, 3
 
Regards
Arun (askIITians forum expert)
Approve this if you find it helpful
Deelip Mukhedkar
6 years ago
Dear student Total no of orbitals =(2l+1)here l=3so if u substitute the value of l u will get answer 7.If u like the solution plz approve it
Ranjith K
5 years ago
Dear student,l=3 indicates that it is an f-subshell. f-subshell always consists of 7 orbitals.Hint:l=0 s-subshelll=1 p-subshelll=2 d-subshell
Anjali kumari
5 years ago
As n= 4 and l= 3. Therefore the number of orbitals=0,1,2,3,-1,-2,-3. That is the number of orbitals= 7.